# Exploiting the Final Component of Generator Architectures for AI-Generated Image Detection

### v1

@@ 4.2 Results & discussions
Q5: How does the classification backbone affect detection performance? To evaluate the impact of the backbone on our detector, we train a model using a pretrained DINOv2 backbone on the VAE training set used in the previous experiments. We compare the performance of DINOv2 and DINOv3 on the testing sets from Q3 and Q4 (Fig. 7). The results show that the newer backbone provides a significant improvement in accuracy. For AP, since the margin between the logits of positive and negative images is sufficiently large with DINOv2, using the newer backbone has little effect.
@@ 5 Conclusion
This study demonstrates that leveraging the final components of modern image generators offers a powerful means to enhance deepfake detection generalization. By contaminating real images through these components and training on minimal representative samples, the proposed detector achieves superior performance across diverse unseen generators, highlighting the value of insights for robust detection.
@@ 6 Discussion on Property 2
For convenience, the paired gradient $\hat{\nabla}_{\theta}^{\text{paired}}$ can be rewritten as:
$\hat{\nabla}_{\theta}^{\text{paired}}=-\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\Big[\nabla_{\theta}\log\theta(\mathbf{x}_{i})+\nabla_{\theta}\log(1-\theta(\varphi(\mathcal{E}(\mathbf{x}_{i}))))\Big].$
@@
$g_{i}=\nabla_{\theta}\log\theta(\mathbf{x}_{i}),\qquad g_{i}^{\prime}=\nabla_{\theta}\log(1-\theta(\varphi(\mathcal{E}(\mathbf{x}_{i})))),$
@@
$G=\{g_{i}\}_{i=1}^{|\mathcal{B}_{t}|},\qquad G^{\prime}=\{g_{i}^{\prime}\}_{i=1}^{|\mathcal{B}_{t}|}.$
Under this notation, $\hat{\nabla}_{\theta}^{\text{paired}}$ becomes
$\hat{\nabla}_{\theta}^{\text{paired}}=-\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\big[g_{i}+g_{i}^{\prime}\big].$
@@
Using the fact that independent random variables have zero covariance, the variance of $\hat{\nabla}_{\theta}^{\text{paired}}$ is:
$\text{Var}(\hat{\nabla}_{\theta}^{\text{paired}})$ $=\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\Big[\text{Var}(g_{i})+\text{Var}(g_{i}^{\prime})+2\text{Cov}(g_{i},g_{i}^{\prime})\Big]$
@@
$\text{Var}(G)=\sum_{i=1}^{|\mathcal{B}_{t}|}\text{Var}(g_{i}),\qquad\text{Var}(G^{\prime})=\sum_{i=1}^{|\mathcal{B}_{t}|}\text{Var}(g_{i}^{\prime}),$
@@
$\text{Var}(\hat{\nabla}_{\theta}^{\text{indep}})=\frac{1}{|\mathcal{B}_{t}|}\text{Var}(G)+\frac{1}{|\mathcal{B}_{t}|}\text{Var}(G^{\prime}).$
@@
$\text{Var}(\hat{\nabla}_{\theta}^{\text{paired}})>\text{Var}(\hat{\nabla}_{\theta}^{\text{indep}}).$
@@
We empirically compare pairwise training with independent training. Using samples generated by VAE_1.1, we train detectors with identical model architectures, hyperparameters, and training data. The only difference lies in how mini-batches are constructed: pairwise training loads matched pairs ($\mathbf{x}$, $\mathbf{x}^{\prime}$) into the same mini-batch, whereas independent training samples $\mathbf{x}$ and $\mathbf{x}^{\prime}$ independently and loads them separately, while ensuring that each mini-batch contains the same number of positive samples $\mathbf{x}$ and negative samples $\mathbf{x}^{\prime}$. Table A reports the results. We copy the baseline numbers and Ours_VAE_1.1 results from Table 2 of the main paper, along with the additional row for pairwise training. The average accuracy of pairwise training is 97.22%, compared to 99.22% for independent training—a 2% gap. This empirical observation suggests that pairwise training does not provide an advantage over independent training, consistent with our theoretical analysis above.
@@
vae.decoder based diffusion.denoiser.super-resolution based vq.decoder based
SD1.3 SD1.4 SD XL Flux SD 3.5 DALL·E 2 DALL·E 3 Glide Emu3 LlamaGen
Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP Acc AP
DIRE (ICCV2023) 48.75 0.43 48.65 0.43 50.00 0.59 61.66 0.73 55.65 0.59 56.45 0.67 50.35 0.54 49.50 0.49 53.65 0.57 54.05 0.59
LGrad (CVPR2023) 54.45 0.46 55.85 0.48 52.60 0.48 81.91 0.92 77.29 0.87 55.65 0.56 47.65 0.40 47.45 0.52 54.80 0.48 60.70 0.63
NPR (CVPR2024) 50.70 0.51 50.60 0.52 52.45 0.56 90.00 0.92 88.04 0.90 49.25 0.50 50.35 0.51 48.80 0.51 51.70 0.53 52.25 0.55
FatFormer (CVPR2024) 52.97 0.54 53.72 0.54 66.67 0.75 55.32 0.90 75.86 0.95 54.12 0.56 35.88 0.34 67.77 0.76 63.72 0.66 77.21 0.87
RINE (ECCV2024) 87.60 0.96 87.15 0.96 74.60 0.86 51.70 0.87 96.49 0.99 85.90 0.94 42.35 0.31 63.95 0.72 62.10 0.70 90.65 0.98
CoDE (ECCV2024) 97.75 0.97 97.30 0.97 75.20 0.75 71.09 0.72 76.35 0.75 54.60 0.54 73.15 0.73 80.75 0.80 77.45 0.77 90.50 0.90
C2P-CLIP (AAAI2025) 75.15 0.93 76.70 0.93 77.70 0.95 53.96 0.97 80.88 0.99 55.60 0.77 63.20 0.90 86.65 0.97 68.70 0.91 83.80 0.98
AIDE (ICLR2025) 57.60 0.60 56.70 0.60 62.70 0.64 98.50 0.99 96.45 0.99 59.75 0.43 14.15 0.31 59.75 0.77 62.50 0.57 63.25 0.85
Ours_vae_1.1 99.70 1.0 99.70 1.0 99.70 1.0 99.17 .9998 99.96 1.0 99.60 1.0 99.70 .9999 98.85 .9992 99.70 .9999 99.70 .9999
Ours_vae_1.1_pairwise 98.20 .9999 98.20 .9999 98.20 .9992 98.88 .9542 99.07 .9775 98.10 .9998 98.20 .9970 98.05 .9991 98.20 .9998 98.20 .9998
@@ 9 Datasets construction
Based on Property 2, we generate the constructed image sets $\mathbf{x}^{\prime}$} offline and save them before training begins. For each of the three representative final components, the corresponding constructed training set is produced using its well-trained, optimized encoder, following the guidance of Property 1. Below, we list the real images and the pretrained $\mathcal{E}^{*}$ and $\varphi^{*}$ used in our experiments:
  • VAE_1.1 training set: The real images are from the full MS-COCO 2014 dataset (118k images). The pretrained VAE used to generate negative samples is the VAE from https://huggingface.co/stabilityai/stable-diffusion-2-1.
  • HiDream training set in Experiment Q1: The pretrained VAE weights are from HiDream-I1-Full https://huggingface.co/HiDream-ai/HiDream-I1-Full.
  • VQ_2.2 training set: The pretrained model used for converting real images is JanusPro-7B https://huggingface.co/deepseek-ai/Janus-Pro-7B. Due to its fixed patch size, Janus only supports a resolution of 378×378. Therefore, we filter out images smaller than this size and center-crop the remaining MS-COCO images to 378×378.
  • Super-resolution_3.3 training set: We use PixelFlow https://huggingface.co/ShoufaChen/PixelFlow-Text2Image/tree/main/model.pt, whose last stage is pretrained to upscale images from 128×128 to 512×512. For real images in MS-COCO, we crop them to 512×512 whenever the resolution allows. We do not downsample and re-upsample the real images; instead, we directly use real images that match the output resolution.
  • Sparse selection set: We run Algorithm 4 on the constructed sample sets from VAE_1.1, VQ_2.2, and Super-resolution_3.3. Once the constructed (negative) samples are selected, their corresponding real images form the positive training set.
For the testing set, we use the full sets from the benchmarks Synthbuster, FakeBench, and WildRF. In Synthbuster, there are 1,000 high-resolution real images from RAISE-1K, along with 1,000 fake images generated by each of nine models: DALL·E2, DALL·E3, Adobe Firefly, MidJourney v5, SD1.3, SD1.4, SD2, SDXL, and Glide. The corresponding columns in Tables 2 and 3 correspond to these subsets. To include more recent generators, we additionally construct the following test sets used in the experiments reported in the main paper:
  • HiDream (Q1), Emu3, and LlamaGen (Q2): Real images are the same as in Synthbuster. Fake images are generated from the same 1,000 prompts used for the original nine generators (listed in prompts.csv in https://zenodo.org/records/10066460/files/synthbuster.zip). Generation parameters and image resolutions follow the defaults in each repository, as listed in Table B.
  • SD3.5 and Flux-1-dev: In the VAE category for Q2, three generators already exist in Synthbuster. We generate additional large-scale test sets from SD3.5 and Flux-1-dev, with 10,000 fake images each. Prompts are short descriptions generated by an LLM. Since corresponding real images are not available, we use the full PASCAL VOC training split, which contains 9,600 real images to compute balanced accuracy for the binary classes.
  • Satellite testing set (Q4): The real images comprise all images from four categories of the fMoW validation set (https://github.com/fMoW/dataset): amusement park (2,556 images) , car dealership (2,156 images), electric substation (2,626 images ), and stadium (2,575 images). Each real image has associated metadata including location (country). Fake images are generated using prompts of the form: “a fMoW satellite image of {class_name} in {country}”. The number of generated images matches the number of real images for each class.

### v2

@@ 5.2 Results across different category of generators
Results Table 1 reports detection results across generator categories. The best results are highlighted in bold, and the best baseline results are indicated with a wavy underline. It is evident that robustness is a key challenge, especially as new generators continue to emerge. DIRE and LGrad, which were trained on GAN-based data, perform poorly on most modern generators. However, for Flux and SD3.5, LGrad achieves 81.91% accuracy and 91.81% AP. Similarly, RINE, benefiting from CLIP-based features, performs well on some generators, such as LlamaGen (90.65% accuracy) and SD3.5 (96.49%). All baselines published before 2025 struggled to detect DALL·E3 images, a challenge that persisted until BFree was introduced. BFree remains the strongest baseline but still shows limited robustness on Glide, Flux, and HiDream generations. Notably, BFree produces high AP scores even when its prediction accuracy is low. For example, it achieves 69.62% / 96.62% (accuracy / AP) on Flux images, and 68.61% / 93.81% on Glide images. This discrepancy indicates that, although the model ranks positive samples effectively, its thresholded predictions may incorrect, which may pose challenges in applications that rely on well-calibrated prediction confidence.
In contrast, our approach performs strongly within each category and generalizes effectively across categories, consistently achieving higher accuracy on various subsets. Remarkably, the detector trained on the sparse set, using only 100 fake samples from each of the three components (300 fake and 300 real images in total), achieves performance comparable to our three models trained on individual components with the full MS-COCO dataset. This small training set provides further evidence of the generalization capability of our approach and offers a practical advantage, as generating large-scale training images is costly.
5.3 Results on wild generations
Testing set (1) Unknown architectures—Firefly and Midjourney subsets from Synthbuster. Both are commercial generators lacking open-source code or public documentation of their architectures or components. (2) Mixed generators—FakeBench[27]. It contains 3,000 real images and 3,000 fake images generated by multiple models. (3) In-the-wild generation—WildRF[6]. It comprises three subsets: Reddit, Facebook, and Twitter, each containing real and synthetic images sourced from social media platforms with unknown origins.
Results Table 3 reports the results. Baseline methods exhibit inconsistent results across benchmarks. For instance, RINE performs well on Firefly (92.10% accuracy) but drops sharply on Midjourney (55.45%) and Facebook (51.56%). BFree performs well on most subsets but degrades on FakeBench. This instability underscores the challenge of real-world generalization, particularly in true zero-shot scenarios where no prior information about the generators is available. In contrast, our detector consistently surpasses all baselines and maintains robust, stable performance across all subsets, demonstrating that leveraging traces from the final component yields strong generalization to diverse architectures and challenging real-world conditions.
5.4 Results on domain-specific fine-tuned generators
Testing set SatelliteDiffusion[21] dataset is generated using a satellite imagery model fine-tuned from SD2 and conditioned on satellite metadata. Real images are sourced from high-resolution satellite imagery, while synthetic counterparts are produced using textual descriptions of the real samples. Four subsets: amusement park, car dealership, electric substation, and stadium are tested, each including around 2,000 real and 2,000 generated images.
Results Table 3 presents the comparison. Our models achieve results closely matching those reported for the original SD2 evaluations in Table 1, demonstrating that the detection approach remains effective for generators fine-tuned on domain-specific images and conditions. Baseline models such as CoDE exhibit a similar phenomenon: on SD2, it achieves 88.05% accuracy, while on fine-tuned subsets such as amusement_park and electric_station, the accuracy slightly decreases to 86.66% and 84.50%, respectively. On the stadium subset, both CoDE and the best baseline, BFree, experience larger drops in performance.
5.5 Ablation study
Different backbones To evaluate the impact of the backbone on our detector, we train an additional model using the pretrained DINOv2 backbone. The comparison of performance between DINOv2 and DINOv3 is shown in Fig. 5. The results show that the newer backbone provides a significant improvement in accuracy. For AP, since the margin between the logits of positive and negative images is sufficiently large with DINOv2, using the newer backbone has little effect.
Final component vs. full pipeline To verify whether traces from the final component alone are sufficient to detect images generated by the full pipeline, we train an additional detector on images constructed from the final component of HiDream-I1 on the MS-COCO dataset using Algorithm 1. To evaluate its performance on full-pipeline outputs, we generate test images following Section 5.2 using the 1,000 Synthbuster prompts. The detector achieves 90.5% accuracy and 99% AP, demonstrating that training on the generator’s final component generalizes effectively to its full outputs.
@@ 6 Conclusion
This study shows that leveraging the final components of modern image generators offers an effective strategy for improving deepfake detection generalization. By contaminating real images through these components and training on minimal representative samples, the proposed detector achieves superior performance across diverse unseen generators, highlighting the value of insights for robust detection.
@@ 1 Discussion on Property 2
For convenience, the paired gradient $\hat{\nabla}_{\theta}^{(P)}$ can be rewritten as:
$\hat{\nabla}_{\theta}^{(P)}=-\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\Big[\nabla_{\theta}\log\theta(\mathbf{x}_{i})+\nabla_{\theta}\log(1-\theta(\varphi(\mathcal{E}(\mathbf{x}_{i}))))\Big].$
@@
$g_{i}=\nabla_{\theta}\log\theta(\mathbf{x}_{i}),\quad g_{i}^{\prime}=\nabla_{\theta}\log(1-\theta(\varphi(\mathcal{E}(\mathbf{x}_{i})))),$
@@
$G=\{g_{i}\}_{i=1}^{|\mathcal{B}_{t}|},\quad G^{\prime}=\{g_{i}^{\prime}\}_{i=1}^{|\mathcal{B}_{t}|}.$
Under this notation, $\hat{\nabla}_{\theta}^{(P)}$ becomes
$\hat{\nabla}_{\theta}^{(P)}=-\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\big[g_{i}+g_{i}^{\prime}\big].$
@@
Using the fact that independent random variables have zero covariance, the variance of $\hat{\nabla}_{\theta}^{(P)}$ is:
$\text{Var}(\hat{\nabla}_{\theta}^{(P)})$ $=\frac{1}{|\mathcal{B}_{t}|}\sum_{i=1}^{|\mathcal{B}_{t}|}\Big[\text{Var}(g_{i})+\text{Var}(g_{i}^{\prime})+2\text{Cov}(g_{i},g_{i}^{\prime})\Big]$
@@
$\text{Var}(G)=\sum_{i=1}^{|\mathcal{B}_{t}|}\text{Var}(g_{i}),\quad\text{Var}(G^{\prime})=\sum_{i=1}^{|\mathcal{B}_{t}|}\text{Var}(g_{i}^{\prime}),$
@@
$\text{Var}(\hat{\nabla}_{\theta}^{(I)})=\frac{1}{|\mathcal{B}_{t}|}\text{Var}(G)+\frac{1}{|\mathcal{B}_{t}|}\text{Var}(G^{\prime}).$
@@
$\text{Var}(\hat{\nabla}_{\theta}^{(P)})>\text{Var}(\hat{\nabla}_{\theta}^{(I)}).$
@@
We empirically compare pairwise training with independent training. Using the training set VAE, we train detectors with identical model architectures, hyperparameters, and training data. The only difference lies in how mini-batches are constructed: pairwise training loads matched pairs ($\mathbf{x}$, $\mathbf{x}^{\prime}$) into the same mini-batch, whereas independent training samples $\mathbf{x}$ and $\mathbf{x}^{\prime}$ independently and loads them separately, while ensuring that each mini-batch contains the same number of positive samples $\mathbf{x}$ and negative samples $\mathbf{x}^{\prime}$. Table A reports the results. We copy the baseline numbers and Ours_vae results from Table 1 of the main paper, along with the additional row for pairwise training. The average accuracy of pairwise training is 97.22%, compared to 99.22% for independent training—a 2% gap. This empirical observation suggests that pairwise training does not provide an advantage over independent training, consistent with our theoretical analysis above.
@@
diffusion.denoiser based vq.de-tokenizer based vae.decoder
DALL·E 2 DALL·E 3 Glide Emu3 LlamaGen HiDream
Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP%
DIRE 56.45 67.05 50.35 53.53 49.50 49.27 53.65 56.73 54.05 59.26 51.40 51.21
LGrad 55.65 55.73 47.65 40.41 47.45 51.83 54.80 48.03 60.70 63.08 40.05 39.06
NPR 49.25 50.18 50.35 50.48 48.80 51.43 51.70 53.05 52.25 55.48 47.75 45.51
FatFormer 54.12 56.02 35.88 33.95 67.77 75.75 63.72 66.10 77.21 87.73 34.48 31.28
RINE 85.90 94.05 42.35 30.83 63.95 72.09 62.10 69.68 90.65 98.36 44.15 39.20
CoDE 54.60 54.26 73.15 72.76 80.75 80.36 77.45 77.06 90.50 90.11 71.45 71.06
C2P-CLIP 55.60 76.87 63.20 90.45 86.65 96.53 68.70 90.60 83.80 97.85 51.20 51.02
AIDE 59.75 42.26 14.15 31.34 59.75 76.71 62.50 56.86 63.25 84.55 48.30 37.95
BFree 93.30 99.16 96.22 99.64 68.61 93.81 99.05 99.97 99.15 99.98 81.40 93.19
Ours_VAE 99.60 100.0 99.70 99.99 98.85 99.92 99.70 99.99 99.70 99.99 99.15 99.94
Ours_VAE_pair 98.10 99.98 98.20 99.70 98.05 99.91 98.20 99.98 98.20 99.98 97.55 99.27
vae.decoder based
SD1.3 SD1.4 SD XL SD2 Flux SD 3.5
Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP% Acc% AP%
DIRE 48.75 43.24 48.65 43.38 50.00 58.90 47.45 40.37 61.66 73.10 55.65 58.00
LGrad 54.45 46.46 55.85 48.14 58.65 53.26 48.75 42.80 81.91 91.81 77.29 87.08
NPR 50.70 51.05 50.60 51.69 52.45 55.54 43.50 46.23 90.00 92.34 88.04 89.82
FatFormer 52.97 53.60 53.72 54.07 66.67 75.43 51.57 54.41 55.32 89.83 75.86 95.39
RINE 87.60 95.96 87.15 95.95 74.60 85.56 70.30 80.74 51.70 86.61 96.49 99.93
CoDE 97.75 97.36 97.30 96.91 75.20 74.81 88.05 87.66 71.09 72.17 76.35 75.26
C2P-CLIP 75.15 92.93 76.70 93.03 77.70 94.67 69.15 92.30 53.96 94.58 80.88 98.87
AIDE 57.60 60.25 56.70 60.14 62.70 63.98 53.45 42.09 98.50 99.80 96.45 99.72
BFree 99.25 99.99 99.15 99.99 99.25 99.98 99.80 99.93 69.62 93.62 98.74 99.94
Ours_VAE 99.70 100.0 99.70 100.0 99.70 100.0 99.70 100.0 99.17 99.98 99.96 100.0
Ours_VAE_pair 98.20 99.99 98.20 99.99 98.20 99.92 98.20 99.93 98.88 95.42 99.07 97.75
@@ 4 Datasets construction
Based on Property 2, we generate the constructed image sets $\mathbf{x}^{\prime}$} offline and save them before training begins. For each of the three representative final components—Stable Diffusion 2.1,JanusPro and PixelFlow, the corresponding constructed training set is produced using its well-trained, optimized encoder, following the guidance of Property 1. Below, we list the real images and the pretrained $\mathcal{E}^{*}$ and $\varphi^{*}$ used in our experiments:
  • VAE training set (Sec 5.1): The real images are from the full MS-COCO 2014 dataset (118k images). The pretrained VAE used to generate negative samples is the VAE from https://huggingface.co/stabilityai/stable-diffusion-2-1.
  • VQ training set (Sec 5.1): The pretrained model used for converting real images is JanusPro-7B https://huggingface.co/deepseek-ai/Janus-Pro-7B. Due to its fixed patch size, Janus only supports a resolution of 378×378. Therefore, we filter out images smaller than this size and center-crop the remaining MS-COCO images to 378×378.
  • Diffusion training set (Sec 5.1): We use PixelFlow https://huggingface.co/ShoufaChen/PixelFlow-Text2Image/tree/main/model.pt, whose last stage is pretrained to upscale images from 128×128 to 512×512. For real images in MS-COCO, we crop them to 512×512 whenever the resolution allows. We do not downsample and re-upsample the real images; instead, we directly use real images that match the output resolution.
  • HiDream training set (Sec 5.5): The pretrained VAE weights are from HiDream-I1-Full https://huggingface.co/HiDream-ai/HiDream-I1-Full.
For the testing set, we use the full sets from the benchmarks Synthbuster, FakeBench, and WildRF. In Synthbuster, there are 1,000 high-resolution real images from RAISE-1K, along with 1,000 fake images generated by each of nine models: DALL·E2, DALL·E3, Adobe Firefly, MidJourney v5, SD1.3, SD1.4, SD2, SDXL, and Glide. The corresponding columns in Tables 1 and 2 correspond to these subsets. To include more recent generators, we additionally construct the following test sets used in the experiments reported in the main paper: